OpenStudy (anonymous):

solve the initial value problem y'+3y=2x+1, y(1)=5

OpenStudy (anonymous):

i know that the integrating factor is e^3x but i'm stuck after that.

OpenStudy (turingtest):

multiply everything through by the integrating factor what do you have after?

OpenStudy (anonymous):

(e^3x)y=(e^3x)(2x-1)

OpenStudy (turingtest):

you skipped a step, which is making you confused as to how to proceed don't 'undo' the product rule on the left without leaving the proper derivative notation what you have written above is not true, until you integrate the right...

OpenStudy (anonymous):

so then its (e^3x)y'+(e^3x)y=2x-1?

OpenStudy (turingtest):

\[y'+3y=2x+1\]\[y'e^{3x}+3ye^{3x}=e^{3x}(2x-1)\]here is the step where we notice the left is a product rule:\[(ye^{3x})'=e^{3x}(2x-1)\]you need to keep y in the parentheses think about it, how could you apply the product rule to the expression as you have it with y' outside the parentheses? now you can integrate both sides...

OpenStudy (anonymous):

I did get to that step its the integrating that i am having trouble with because i know its integration by parts and i got the RHS = {(1/3)e^3x}(4/3 + 2x)+c it just doesnt seem right to me

OpenStudy (turingtest):

if that's not right it's very close I'm trying to see if our answers simplify the same I will do the integration...

OpenStudy (turingtest):

\[\int(2x-1)e^{3x}dx=\frac13(2x-1)e^{3x}-\frac23\int e^{3x}dx\]\[=\frac13(2x-1)e^{3x}-\frac29e^{3x}+C=(\frac23x-\frac59)e^{3x}+C\]I think you're off by a 1 somewhere...

OpenStudy (anonymous):

the equation starts with +1 not -1

OpenStudy (turingtest):

ah so it does... minor change, one sec.

OpenStudy (turingtest):

\[\int(2x+1)e^{3x}dx=\frac13(2x+1)e^{3x}-\frac23\int e^{3x}dx\]\[=\frac13(2x+1)e^{3x}-\frac29e^{3x}+C=\frac19(6x+1)e^{3x}+C\]there we go :)

OpenStudy (anonymous):

so now do i just divide by the integrating factor?

OpenStudy (turingtest):

yes, then apply the initial conditions to find C

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (turingtest):

very welcome :D

OpenStudy (anonymous):

i got c=45/7 ?

OpenStudy (turingtest):

no, there should be an e in it... what do you have after dividing by the IF ?

OpenStudy (anonymous):

i did 5=[(1/9)(6(1)+1)3^(3*1)]/e^3(1) + c/e^3

OpenStudy (turingtest):

you need to be more careful, one step at a time also, when dividing by e^(something) it's much nicer to just multiply both sides by e^(-something) makes the expression much simpler-looking....

OpenStudy (turingtest):

we had\[(ye^{3x})'=e^{3x}(2x+1)\]\[ye^{3x}=\frac19(6x+1)e^{3x}+C\]now multiply both sides by e^(-3):\[y=\frac19(6x+1)+Ce^{-3x}\]you forgot to cancel out that other e^(3x) earlier... now initial conditions...

OpenStudy (turingtest):

\[y(1)=\frac19(6+1)+Ce^{-3}=5\]\[Ce^{-3}=5-\frac79=\frac{38}9\to C=\frac{38}9e^3\]so we get\[y=\frac19(6x+1)+\frac{38}9e^3e^{-3x}\]and just to make it prettier we can simplify\[y=\frac19(6x+38e^{3(1-x)}+1)\]

OpenStudy (turingtest):

*slight typo in the post before last: after\[ye^{3x}=\frac19(6x+1)e^{3x}+C\]we multiply both sides by e^(-3x) to get... (the rest is correct)

OpenStudy (anonymous):

THANK YOU

OpenStudy (turingtest):

you're welcome here are some good notes on the subject: http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx

OpenStudy (turingtest):

I'm not really a stickler about it, but it's customary to give medals for good explanations so others know that the problem has been solved. So feel free to click 'good answer' if you are satisfied with my explanation. welcome to Open Study!